OpenStudy (anonymous):

A car travels at a uniform acceleration over a distance of 440.0 m. Its final velocity after 20.0 s is 33.0 m/s. What is its initial velocity? 0.0 m/s, 11.0 m/s, 22.0 m/s, or 8.7 m/s **Not sure how this one goes! :( thank you!

OpenStudy (abhisar):

WHat equation will u use here ?

OpenStudy (anonymous):

it's one of the ones you wrote earlier right?

OpenStudy (anonymous):

is it the v=u+gt?

OpenStudy (abhisar):

Yes....what quantities are given in the question ?

OpenStudy (abhisar):

Time is not given !

OpenStudy (anonymous):

initial velocity=0 final velocity=33 distance=440 time=20 s ?

OpenStudy (anonymous):

oh so would it be 0.0 m/s ? or not sure if i got those values right? :/

OpenStudy (abhisar):

How time is 20s ? Recheck

OpenStudy (anonymous):

not sure :/ what does it mean when it says "final velocity after 20.0 s" ? that's not the total time then? and @jim_thompson5910 which one applies here? :/ not sure, but i really need to memorize all of those!!

jimthompson5910 (jim_thompson5910):

you need to find what piece of info you don't have (and that you aren't solving for) then use the chart to locate the X that corresponds to that variable to get the correct formula

OpenStudy (anonymous):

okay:) so vf=final velocity=33 m/s ? would that be the final? not sure:(

jimthompson5910 (jim_thompson5910):

"A car travels at a uniform acceleration over a distance of 440.0 m. Its final velocity after 20.0 m is 33.0 m/s. What is its initial velocity?" you want to find vi you aren't given the time t, so look for the formula with an X under t to get the formula you need

OpenStudy (nincompoop):

use \(\sf \large s = \frac{1}{2} (v_f-v_i) \times t\) then solve for what you're looking for

jimthompson5910 (jim_thompson5910):

oh is that 20 s and not 20 m?

jimthompson5910 (jim_thompson5910):

I see now

OpenStudy (abhisar):

Yes i think so..its 20s

jimthompson5910 (jim_thompson5910):

you aren't given the acceleration and you don't need to solve for that so look for an X under the acceleration 'a' to get your formula

OpenStudy (anonymous):

oops my bad :P typed in m instead of s sorry you guys!!!

OpenStudy (anonymous):

so d=1/2(vf+vi)t ?

OpenStudy (abhisar):

Now u'll hv to calculate a first..

OpenStudy (anonymous):

d=1/2(vf+vi)t d=distance=440 vf=33 vi=? t=20 so 440=1/2(33+x)20 440=1/2(660+20x) 440=330+10x 110=10x x=11 ?

OpenStudy (nincompoop):

ye boy

OpenStudy (abhisar):

u'll have to use 2nd equation to calculate a first. Then u can calculate u from first equation.

jimthompson5910 (jim_thompson5910):

11 m/s looks good

OpenStudy (anonymous):

haha woohoo!! yay! :) thanks so much!! and sorry again for my typo!! will triple check from now on! hehe :)

OpenStudy (nincompoop):

that is your initial velocity, boy!

OpenStudy (nincompoop):

do you want to learn how to derive that or what?

OpenStudy (raffle_snaffle):

i'd like to know how to derive that nin

OpenStudy (nincompoop):

memorizing formula is easy, but knowing how to get the formula is better

OpenStudy (nincompoop):

fifty bucks, @raffle_snaffle I already helped you in modern physics for free :P

OpenStudy (raffle_snaffle):

LOL fifty bucks?